OpenStudy (anonymous):

Solve for x in the population: x over x+12 = 1 over x-1 A. x ≈ −6.02 and x ≈ 5.02 B. x ≈ −3.46 and x ≈ 3.46 C. x ≈ −3.14 and x ≈ 4.14 D. x ≈ −2.61 and x ≈ 4.61

OpenStudy (anonymous):

\[\frac{x}{x+12}=\frac{1}{x-1}\]?

OpenStudy (anonymous):

"cross multiply" get \[x(x-1)=x+12\] \[x^2-x=x+12\] then solve the quadratic equation \[x^2-2x=12\] can you finish it?

OpenStudy (anonymous):

umm no...I wanted to be shown how to get the answer.

OpenStudy (anonymous):

ok you have a choice write \[x^2-2x-12=0\] and use the quadratic formula or keep it at \[x^2-2x=12\] and complete the square this one does not factor using whole numbers so you cannot do that

OpenStudy (anonymous):

since the coefficient \(-2\) it is easiest to complete the square start with \[x^2-2x=12\] half of 2 is 1 and 1 squared is also 1 so write \[(x-1)^2=12+1=13\] take the square root, don't forget the \(\pm\) get \[x-1=\pm\sqrt{13}\] solve for \(x\) get \[x=1\pm\sqrt{13}\] to get a decimal use a calculator for \(1+\sqrt{13}\) and \(1-\sqrt{13}\)

OpenStudy (anonymous):

hmmm okay im ganna see

OpenStudy (anonymous):

No I need the answer...sorry